OpenStudy (anonymous):

integral of [e^x]^2 is ? any help would be great

OpenStudy (tkhunny):

There might be a simplification. Have you considered your exponent Rules?

OpenStudy (perl):

$$ \large \int (e^{x})^2 ~dx \\ = \large \int e^{x}*e^x ~dx\\ =\large \int e^{x+x} ~dx\\ $$ can you solve it now?

OpenStudy (anonymous):

OHH. It would be e^2x then

OpenStudy (perl):

$$ \large \int (e^{x})^2 ~dx \\ = \large \int e^{x}\cdot e^x ~dx\\ =\large \int e^{x+x} ~dx\\ $$

OpenStudy (perl):

yes :)